1. a triangle is 24 feet , 7 yards, 132 inches and what is the perimeter ? & 2. the sum of four consective odd numbers is 216 . Find all four numbers . (I NEED THE HELP NOW )
Answer by ankor@dixie-net.com(22740)

1. A triangle is 24 feet, 7 yards, 132 inches and what is the perimeter ? &
Change to feet
7 * 3 = 21 ft
132/12 = 11 ft
p = 24 + 21 + 11
p = 36 ft or 12 yds
:
2. the sum of four consecutive odd numbers is 216 .
n + (n+2) + (n+4) + (n+6) = 216
4n + 12 = 216
4n = 216 - 12
4n = 204
n = 204/4
n = 51 is the 1st number
:
Find all four numbers; I'll let you find the other 3 numbers